OpenStudy (anonymous):

Checking My Answers What is the slope of the line that passes through the pair of points? (–2, 7), (18, 1) I got -9/4

OpenStudy (anonymous):

it's wrong. answer is -3/10

OpenStudy (anonymous):

Answer is 3/11. Not what the person wrote above me. Slope is rise/run, change in y, over change in x. 1 minus a negative is a positive, and 18 minus 7 is 11.

OpenStudy (anonymous):

@kelberry you're a stupid and you're wrong. firstly learn, then teach. ok??

OpenStudy (anonymous):

I'm not incorrect from my understanding, I'd advise you to look over your answer. If you believe I'm wrong, please show your work to prove it.

OpenStudy (anonymous):

slope=(y2-y1)/(x2-x1) slope=(1-7)/(18-(-2)) slope=-6/20=-3/10 do you understand foolish boy??